OpenStudy (anonymous):

Look at this expression. a3b3c a-3b-3c-1 Which of the following is the simplest form of the expression above?

OpenStudy (aravindg):

a-3b-3c-1 and a3b3c are two different expressions ?

OpenStudy (anonymous):

A. -1 B. 1 C. a-6 b-6c-2 D. a6 b 6c2

OpenStudy (anonymous):

I am not totally Sure on How to do this problem. Are you suppose to add The A To the A and B to the B And C to The C ?

OpenStudy (anonymous):

\[\huge \color{green}{\textbf{Welcome To Openstudy...}}\]

terenzreignz (terenzreignz):

Is it a fraction?

OpenStudy (anonymous):

a3 b3 c --------- a-3 b-3 c-1 That is how the problem is written

OpenStudy (anonymous):

It looks fine now..

OpenStudy (aravindg):

ok now I understand a^-1=1/a^3 It will be similiar to b and c in denomiator

OpenStudy (anonymous):

\[\large \frac{a^3 \cdot b^3 \cdot c}{a^{-3} \cdot b^{-3} \cdot c^{-1}}\]

OpenStudy (anonymous):

\[\frac{x^a}{x^b} = x^{a - b}\]

OpenStudy (aravindg):

so you are left with